OpenStudy (millsemily):

Nicole is factoring 0.49t^18 - 25 by using the rule a^2 - b^2 = (a + b)(a - b). What will she use for the value of a? A. 0.7t^9 B. 0.07t^9 C. 0.49 D. 0.7 I'd love some help with this please :)

OpenStudy (jackthegreatest):

lol thats somewat hard

OpenStudy (jackthegreatest):

i thought u were talking about factoring stuff like 4

OpenStudy (jackthegreatest):

but i think the answer is a

OpenStudy (jackthegreatest):

dont take my answer its a guess

OpenStudy (millsemily):

Told you :)

OpenStudy (millsemily):

Haha, it's alright :)

OpenStudy (bibby):

if \(0.49t^{18} = a^2\), and \(\large{a^b}^c=a^{b*c}\) you can rewrite it as \(0.7^2*{a^9}^2\)

OpenStudy (bibby):

and so \(a=0.7a^9\)

OpenStudy (millsemily):

I thought it was B

OpenStudy (bibby):

ugh by a I meant t

OpenStudy (bibby):

\(0.49t^{18} = a^2\) \(a=0.7t^9\)

OpenStudy (jackthegreatest):

wait emily those carrots mean "exponents"?

OpenStudy (jackthegreatest):

so my guess was correct @bibby ?

OpenStudy (millsemily):

Wait. So that means the answer is C? Right

OpenStudy (millsemily):

Yes Jack :) ^= exponents

OpenStudy (jackthegreatest):

ahhh

OpenStudy (jackthegreatest):

then i may be able to help

OpenStudy (millsemily):

It's A. If you plug it in, I got that answer.

OpenStudy (jackthegreatest):

i think it is actulaly a

OpenStudy (millsemily):

It's okay Jack! Bibby showed me the right thing. I confused what was where.

OpenStudy (millsemily):

I got it now :)

OpenStudy (bibby):

no the answer would be a 0.49t^18 - 25 by using the rule a^2 - b^2 = (a + b)(a - b). if we're treating \(0.49t^{18}\) as \(a^2\) and we want a, we can take the sqrt of both parts because \((ab)^m=a^mb^m\)

OpenStudy (bibby):

so the sqrt of 0.49 is 0.7 and \(\sqrt{t^{18}}={t^{18}}^\frac{1}{2}=t^{\frac{18}{2}}=t^9\)

OpenStudy (jackthegreatest):

so a is correct?

OpenStudy (bibby):

sorry my brain is loopy

OpenStudy (jackthegreatest):

can u show me how to do it @bibby i want to learn ahead

OpenStudy (millsemily):

Yeah I got that. A is correct. Thank you :)

OpenStudy (millsemily):

Haha better then mine bibby